OpenStudy (stormswan):

Please help. I will give a medal and fan. Solve each system using substitution. Check your solution. m=5p+8 m=-10p+3

OpenStudy (anonymous):

It gives you two equations for the variable m.. So it depends on your preference of either substituting (-10p+3) for m into the first equation or substituting (5p+8) into the second equation.. So which one do you prefer?

OpenStudy (stormswan):

im not so sure what that means, given i don't know any of this stuff, but lets go for the second one?

OpenStudy (anonymous):

Okay, let's start from the beginning if you don't know how to do it. Are you saying that you don't know how to do substitution?

OpenStudy (stormswan):

yeah.

OpenStudy (anonymous):

Here is a website that'll help you... http://www.mathsisfun.com/algebra/substitution.html read this first and then come back :)

OpenStudy (stormswan):

hmm, i knew that... but I meant like i dont understand how to solve the problem and stuff.

OpenStudy (anonymous):

Okay. You have two equations for the variable m. Both equations EQUAL m.

OpenStudy (anonymous):

m is equal to both (5p+8) and (-10p+3).

OpenStudy (anonymous):

So if m is equal to both those two equations.. Those two equations are equal to each other. It's like saying a = b = c.

OpenStudy (anonymous):

So what can you do with those two equations to solve for x?

OpenStudy (stormswan):

you can add them together?

OpenStudy (anonymous):

If a = b, it wouldn't be the same thing as a + b , right? For example. 2=2 is not the same as 2+2.

OpenStudy (stormswan):

wait... you meant to solve m? not x?^^

OpenStudy (anonymous):

Yes, sorry, m. :)

OpenStudy (stormswan):

ok, and yeah. thats not the same thing. so what do we do?

OpenStudy (anonymous):

If a=b, and b=c.... You could substitute c in for b, so a=c. Do you get that part? :P

OpenStudy (stormswan):

but then where would the b go? lol

OpenStudy (anonymous):

That's the point of substituting. You're substituting so that you can get rid of the b.

OpenStudy (stormswan):

oh, ok :) i get it.

OpenStudy (stormswan):

but in this case, there isnt a third letter..

OpenStudy (anonymous):

Yes, that was just an algebraic law that I was explaining to you. The name's slipping my mind right now but it's an accepted rule.

OpenStudy (anonymous):

Okay, now that you get that part, we can look at the two equations again. m=5x+8, m=-10x+3.

OpenStudy (stormswan):

yup

OpenStudy (anonymous):

Oops, p, not x. m=5p+8, m=-10p+3. If both (5p+8) and (-10p+3) are equal to m, what can you do? BOTH of them are equal to m.

OpenStudy (anonymous):

This is like saying an apple is a fruit, and a fruit is a food. Therefore, an apple is a food.

OpenStudy (stormswan):

well, you can make them equal then? since you cant add them?

OpenStudy (anonymous):

YES, bingo! You set them equal to each other because they're BOTH equal to m, which makes them equal.

OpenStudy (anonymous):

Now, if you set them equal to each other, what would you get for the value of p?

OpenStudy (stormswan):

im not sure how to make them equal though haha thats the problem.

OpenStudy (anonymous):

(5p+8) is equal to (-10p+3), right?

OpenStudy (anonymous):

the = sign basically means "is equal to"

OpenStudy (anonymous):

Do you get it now? :)

OpenStudy (stormswan):

well, yeah. but how do I make them into 1 equation?....

OpenStudy (stormswan):

@dixiemitsy I mean isnt that what I'm supposed to do?

OpenStudy (anonymous):

Sorry, my internet got disconnected! And yes... if you set them equal to each other, 5p+8=-10p+3. Now you can just collect like terms. Put the p's on one side and the numbers on the other. You can then solve for p.